OpenStudy (anonymous):

Solve: -3sqrt(x+7)+10= -1 for x

OpenStudy (anonymous):

is x plus 7 plus 10 all in a square root?

OpenStudy (anonymous):

no, just the x+7

OpenStudy (anonymous):

okay.are they looking for a decimal answer or an answer that allows a square root?

OpenStudy (anonymous):

just the x no square root i think :)

OpenStudy (anonymous):

oh i see i see. i thought it said x = -1

hero (hero):

Graphic girl, I answered your question

OpenStudy (anonymous):

Try to check it again Hero :)

hero (hero):

Graphic girl, what makes you think that F = 3?

hero (hero):

You can't just divide 6/2 The 2 is being raised to a power

OpenStudy (anonymous):

I know. we're supposed to simplify. like make it 6= 2 x 3 over 2 x 2^6

hero (hero):

The only way to re-write it is like this: \[6 \dot\ 2^{-7}\]

OpenStudy (anonymous):

giaxlove. x should be about 6.4

OpenStudy (anonymous):

thnxs :)

OpenStudy (anonymous):

ooh that makes sense but i don't know why my online program says it's 3 times....etc etc

OpenStudy (anonymous):

ohhhh i know whaat it is!

hero (hero):

If you did it that way then it would be 3*2^-6

OpenStudy (anonymous):

it's 3 times 2^-6

OpenStudy (anonymous):

mmm.... graphic girl its wrong :(

hero (hero):

I think my original answer is most correct though. Your online program failed

OpenStudy (anonymous):

thnxs for trying anyways

OpenStudy (anonymous):

haha no cuz it wanted it simplifieeed. and i got it right! yay thanks for your help hero :). giax love---i was wrong? what is the correct answer supposed to be?

hero (hero):

Simplification step was unnecessary

OpenStudy (anonymous):

*shrugs* thats what it asked for heh. but yea i completely see your view too.

hero (hero):

Glad I was able to help you

OpenStudy (anonymous):

ugh won't let me post! but um yeah idk the answer, which is why i'm here lol asking for help :)

OpenStudy (anonymous):

giaxlove. this is how i solved it. I subtracted the 10 from both sides. then i squared both sides. then i simplified

OpenStudy (anonymous):

ok so i got to the part where its sqrt(x=7)==11/3 then i'm stuck

OpenStudy (anonymous):

so i went to -3sqrt(x+7) +10= -1 to -3sqrt(x+7) = -11 to 9(x+7) = 121 and then to 9x +63 = 121 and then...you know the rest

OpenStudy (anonymous):

huh? whered you get the 9?

OpenStudy (anonymous):

she got the 9 by squaring both sides of the equation